Fermi National Laboratory


Accelerator Update for August 10, 2005 - August 12, 2005

Wednesday, August 10
The day shift began with Operations monitoring store 4322 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Tevatron (TeV) experts began conducting end-of-store studies at 8:45 AM.

Operations terminated store 4322 at 9:17 AM.

Operators prepared CDF and D0 for access into their collision halls at 9:42 AM.

Cryo system technicians reported at 9:46 AM that they were beginning to replace the flywheels on TeV sectors B3 and E3 wet engines.

A Main Injector (MI) expert began a slipstacking study at 11:17 AM.

The CDF and D0 accesses were completed by 2:18 PM.

Operations put the TeV through a dry squeeze at 2:36 PM.

Operations put the TeV into shot setup at 3:03 PM.

Operations established combination store 4324 at 6:19 PM with an initial luminosity of 86.95E30. Pbar resumed stacking.

At 7:47 PM, Operations reported that a Linac RF station (KRF2) tripped off for the first of four times in ten minutes. An expert investigated and, after making adjustments, had the station operating normally by 8:01 PM.

At 8:05 PM, Operations reported that NuMI was tripping off whenever a MI RF station (MIRF2) tripped off. Operators contacted an expert who suggested bypassing the station.

Operations began preparing the Recycler for shot setup at 10:40 PM. At 11:58 PM, Operations reported beam problems that they needed expert help with. The shot setup was put on hold.

Thursday, August 11
The midnight shift began with Operations monitoring store 4324 and the Recycler stash, with Pbar and the Recycler stashing antiprotons, with D0 in access to replace a power supply, with MiniBooNE and NuMI taking beam, and with MCenter off due to shot setup.

The Recycler expert helped conduct a successful pilot transfer of antiprotons at 12:11 AM.

The stash ended and Pbar resumed stacking at 12:33 AM.

Operations established beam to MCenter at 12:46 AM.

At 2:29 AM, NuMI tripped off for the first of many times over the next five hours due to beam problems.

The Recycler began stashing antiprotons around 4:24 AM. The stash was completed and Pbar resumed stacking at 5:36 AM.

At 8:11 AM, a TeV antiproton abort kicker (C:AAK5) tripped off and wouldn't reset. Experts thought that a high voltage power supply had failed and would have to be replaced.

Operations established beam to MTest at 9:06 AM.

At 9:47 AM, LRF3 tripped off for the first of seven times in twenty minutes due to the tube arcing. After the last trip experts made adjustments.

At 11:06 AM, LRF3 tripped off four times in fifty minutes.

CDF and D0 turned off parts of their detectors while TeV experts tried to turn on C:AAK5.

At 1:04 PM, Operations reported that a Booster kicker (MKS05) tripped off. Experts reported they completed reterminating a cable at 1:46 PM.

TeV experts reported that the AAK5 kicker problem wasn't due to a failed power supply, but to arcing. This, they said, would take eight hours or more to repair and then require a long conditioning time. The current trouble was how to remove the antiprotons with this kicker off. The experts devised a way of slowly getting rid of the particles without causing a TeV quench. They began this process at 2:21 PM, which also marked the time that store 4324 ended.

TeV experts reported at 3:06 PM that they found a failed filter capacitor for AAK5. This means the repair will take less time to complete.

At 3:13 PM, MI RF tripped off due to a LCW pressure drop. Experts investigated and discovered a LCW leak on MIRF6. Experts had the leak repaired by 4:01 PM.

Operations reported at 4:30 PM that TeV experts had successfully removed the antiprotons from the TeV.

Operations prepared D0 for access at 4:34 PM.

TeV experts reported at 5:18 PM that they had completed repairs on AAK5 and were beginning to condition the kickers.

MTest experimenters said at 8:51 PM that they were done for the night.

Operations reported at 8:54 PM that the A0 kicker systems were ready for running.

Operations put the TeV into a dry squeeze at 9:01 PM.

Operations put the TeV into shot setup at 10:19 PM.

Friday, August 12
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

At 12:03 AM, Operations reported that the fourth transfer of antiprotons, which was to come from the Recycler, wasn't extracted. At 12:22 AM, another Recycler transfer was lost. It appeared that the Recycler RF had collapsed. Operators investigated, contacted and expert, reestablished the RF structure, and completed the transfers.

At 1:36 AM, MIRF2 tripped off for the first of many times right after it was reset. Operators bypassed the station.

Operations established combination store 4328 at 1:44 AM with an initial luminosity of 99.8E30. Pbar resumed stacking and MCenter resumed taking beam.

At 2:48 AM, NuMI tripped off for the first of many times for the rest of the shift due to the MIRF. MI had two stations down. Later in the shift, operators lowered MI's intensity to help solve this problem.

Recycler began stashing antiprotons at 5:32 AM. They completed their stash and Pbar resumed stacking at 6:29 AM.

The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/12/2005   email Fermilab